# Trainable Dynamic Mask Sparse Attention

### v1

@@ Abstract
In large language models, the demand for modeling long contexts is constantly increasing, but the quadratic complexity of the standard self-attention mechanism often becomes a bottleneck. Although existing sparse attention mechanisms have improved efficiency, they may still encounter issues such as static patterns or information loss. We introduce a trainable dynamic mask sparse attention mechanism, Dynamic Mask Attention, which effectively utilizes content-aware and position-aware sparsity. DMA achieves this through two key innovations: First, it dynamically generates content-aware sparse masks from value representations, enabling the model to identify and focus on critical information adaptively. Second, it implements position-aware sparse attention computation that effectively skips unnecessary calculation regions. This dual-sparsity design allows the model to significantly reduce the computational complexity of important information while retaining complete information, achieving an excellent balance between information fidelity and computational efficiency. We have verified the performance of DMA through comprehensive experiments. Comparative studies show that DMA outperforms multi-head attention, sliding window attention, multi-head latent attention, and native sparse attention in terms of perplexity under Chinchilla Scaling Law settings. Moreover, in challenging multi-query associative recall tasks, DMA also demonstrates superior performance and efficiency compared to these methods. Crucially, in the evaluation of a 1.7B parameter model, DMA significantly outperforms multi-head attention in both standard benchmark performance and the challenging needle-in-a-haystack task. These experimental results highlight its capability to balance model efficiency and long-context modeling ability effectively. Our computational kernel code is currently open-sourced at https://github.com/SmallDoges/flash-dmattn, facilitating further research and applications in the community.
@@ 1 Introduction
Recent advances in large language models (LLMs) have enabled remarkable achievements in tasks that require reasoning over long contexts [46], such as deep reasoning [27], codebase generation [58], and multi-turn autonomous agents [42]. A key factor behind these successes is the effective modeling of long-range dependencies, often spanning thousands of tokens [48, 14, 23]. However, the standard self-attention mechanism [50] employed by Transformer architectures inherently suffers from quadratic computational complexity [55], significantly restricting scalability to longer sequences. Consequently, designing attention mechanisms that achieve computational efficiency without sacrificing modeling accuracy has become a crucial research frontier in scaling LLM capabilities.
Existing Methods. Current efficient attention strategies mainly leverage two types of sparsity: take advantage of sparsity of softmax attention [36] for efficient computation of the necessary query-key pairs, and utilize sparsity of long content [21] for selective computation of relevant tokens. These strategies can reduce computational overhead while maintaining performance. Methods that effectively utilize attention sparsity include sliding window attention [4] and native sparse attention [54]. Although these methods can achieve efficient long-context modeling, they struggle to maintain effectiveness. On the other hand, methods that effectively utilize the sparsity of chain-of-thought reasoning include KV cache eviction methods ([31], [59], [60]), block KV cache selection methods ([20], [47], [52]), and filtering methods based on sampling [8], clustering [33], or hashing [15]. These methods discard KV pairs predicted to be irrelevant so that the active cache stays small. However, they often fail to achieve their theoretical speedups in practical deployments due to overhead in dynamic computations or inaccurate sparsification decisions.
Key Challenges. To address these limitations, efficient and effective sparse attention mechanisms must simultaneously address two key challenges: utilizing self-attention sparsity for necessary computation [9] and utilizing long-context sparsity for selective computation [11]. These two requirements are crucial for achieving efficient and effective long-context inference and training in practical applications. Considering both aspects, existing methods still have limitations, often facing fundamental trade-offs between efficiency and effectiveness. For example, modern language modeling has two mainstream models: state space models [22] achieve linear computational complexity by compressing historical information into a fixed-size state, but struggle to recall specific information from the distant past accurately; while Transformers [50] maintain a complete key-value cache for all past tokens, allowing precise retrieval of dependencies between any token pairs, but at the cost of $O(n^{2})$ computational and memory complexity. This contradiction underscores the pressing need to develop attention mechanisms that preserve information integrity while achieving computational efficiency.
Figure 1: Dynamic Mask Attention Architecture. Left: Content-Aware Mask Computation. The mask computation part of dynamic mask attention. In the outer loop, the stride weight $\Delta$ and gate weight $A$ are loaded into high-speed SRAM, and in the inner loop, the zero-order hold method is used to loop through the $V$ blocks loaded into SRAM, sampling from it to generate content-aware $K$ masks. These masks are then causally broadcast to the length of $Q$ in HBM. Finally, in the outer loop, all mask blocks are concatenated to form the final content-aware sparse dynamic mask. Right: Position-Aware Weight Computation. The weight computation part of dynamic mask attention, where in the outer loop, the $K$ and $V$ blocks are looped and loaded into SRAM, and in the inner loop, the $Q$ blocks are accessed, loaded into SRAM, and the output of the attention weight computation is written back to HBM. If the current position of the $K$ block is designated as masked in the dynamic mask, the attention weight at that position is directly filled with 0, skipping the computation at that position, forming the final position-aware sparse attention weights.
@@
Table 1: Comparison of Different Attention Variants. Comparison of different Self-Attention mechanisms. $n$ denotes sequence length, $d_{h}$ represents head dimension, $w$ is window size, $d_{c}$ is compressed dimension, $B$ is compression block size, and $k$ is the number of selected blocks. Complexities focus on attention weight computation and memory requirements.
Mechanism Comp. Complexity Mem. Complexity Mask Sparsity Weight Sparsity
Multi Head Attention $O(n^{2}d_{h})$ $O(n^{2})$ Static Dense
Sliding Window Attention $O(nwd_{h})$ $O(nw)$ Position-aware Position-aware
Multi Head Latent Attention $O(n^{2}d_{c})$ $O(n^{2})$ Static Low-rank Approx
Native Sparse Attention $O(n^{2}d_{c}/B+nkBd_{h}+nwd_{h})$ $O(n^{2}/B+nkB)$ Hybrid Content-adaptive
Dynamic Mask Attention $O(nwd_{h})$ $O(nw)$ Content-aware Position-aware
Method. In order to achieve efficient and effective sparse attention mechanisms, we ingeniously integrate the strengths of State-Space [22] and Self-Attention [50], attempting to strike a balance between the two, and propose Dynamic Mask Attention to address the challenges of long context modeling. As shown in Figure 1, DMA is a trainable content-position dual-aware sparse attention mechanism that retains a complete KV cache to ensure original fidelity and precise recall potential of historical information while avoiding information bottlenecks from fixed-state compression, and learns to identify and focus only on the most relevant segments of the V cache for the current input, enabling dynamic attention mask adjustment that combines position-aware and content-aware sparse computation to allow the model to focus on relevant tokens while ignoring irrelevant ones. At the same time, we have designed dedicated computational kernels for dynamic mask attention, enabling hardware-level skipping of computations in masked regions without additional invalid computations. This allows dynamic mask attention to address the two core innovations that meet the above key requirements: handling self-attention with position-aware sparsity and handling long-context with content-aware sparsity. This optimization enables dynamic mask attention to support efficient and effective inference and training with long context. This optimization enables dynamic mask attention to support efficient and effective inference and training with long context. Table 1 provides a comparative overview of DMA compared to other mainstream attention variants, highlighting its unique characteristics in terms of complexity and sparsity.
Contributions. We comprehensively evaluate the efficiency and effectiveness of Dynamic Mask Attention across multiple dimensions. We compare the performance of different attention variants with the same number of parameters in terms of pre-training perplexity, the challenge version of multi-query associative recall tasks [2], and the speed of completing multi-query associative recall tasks. We also assess the acceleration effects of different implementation kernels in various application scenarios. Additionally, we pre-train DMA and vanilla self-attention on a 1.7B Transformer backbone with 40B tokens and compare their performance on downstream benchmarks, including the needle-in-a-haystack task [29]. Experimental results show that DMA achieves better performance than vanilla attention while outperforming existing efficient sparse attention methods. These results validate that our learnable dynamic mask sparse attention design effectively balances model efficiency and effectiveness. Our computational kernel code is open-sourced at https://github.com/SmallDoges/flash-dmattn to facilitate further research in the community.
2 Related Work
Attention Variants. The landscape of attention mechanisms has evolved since the introduction of the Transformer. Original Multi-Head Attention, while enhancing model representational capacity by parallelizing self-attention operations across different subspaces, suffers from $O(n^{2})$ complexity, performing undifferentiated computation across all token pairs, failing to leverage inherent sparsity, and lacking selective computation driven by chain-of-thought. To mitigate costs, Sliding Window Attention [4] reduced complexity to $O(n\cdot w)$ by limiting attention locally, but still computed densely within windows, and its static mask hindered attending to crucial out-of-window information, detrimental to long-range chain-of-thought. Another approach, Multi-Head Latent Attention [32] used low-rank decomposition to cut complexity and parameters, but at the cost of fine-grained information; this global compression isn’t a chain-of-thought driven selective strategy and can disrupt relative position encoding. On the hardware side, Native Sparse Attention [54] designed hardware-aligned regularized sparse patterns, achieving high computational efficiency, but its static patterns cannot dynamically adjust based on input content, making it challenging to attain accurate selective computation to adapt to the diversity of chain-of-thought. These variants attempt to address efficiency issues but largely fail to achieve dynamic, content-aware sparsity and selective computation.
KV Cache Selection. Another class of methods for long-context efficiency optimizes the KV cache, primarily reducing memory by selectively retaining key-value pairs rather than directly optimizing the sparsity of attention computation. For instance, methods like SnapKV [31], H2O [59] and DAM [57] attempt to retain important tokens, but their criteria, such as historical patterns or recency, often do not align with the selective computation needs of the current chain-of-thought, and attention computation among filtered tokens may still be dense. Coarser-grained selection mechanisms, such as SeerAttention [20] for selecting cache blocks and Quest [47] for selecting cache pages, make it even more challenging to meet the fine-grained selective computation required by chain-of-thought sparsity precisely. Additionally, methods based on sampling or hashing, like MagicPIG [8] and HashAttention [15], reduce the number of KV pairs to attend to through approximation or lookup, but they do not address the problem of achieving content-based dynamic sparsification at the attention computation level.
@@ 3 Background
Figure 2: Sparsity in Language Modeling Tasks. The tasks of Copy [44], Select [1], and Induce [40] are three essential tasks for language modeling. The Copy task requires maintaining a fixed distance between input elements and output elements, the Select task involves selectively remembering or ignoring certain elements based on the input, and the Induce task requires retrieving answers through associative recall based on context. Where the colored parts represent the tokens that the model needs to remember in the current time step of inference, the black parts represent the output tokens that the model needs to predict based on the input, and the white parts represent irrelevant tokens that can be filtered out.
@@ 3.1 Language Modeling Tasks
Sparsity in Language Modeling. As shown in Figure 2, long-context language modeling involves three fundamental tasks: Copying, Selecting, and Inducing. The Copy task requires the model to maintain a fixed-distance relationship between the input and output; the Select task requires the model to remember or ignore specific elements based on their content selectively; and the Induce task requires the model to retrieve answers through associative recall based on context. These three tasks naturally exhibit different sparsity patterns: positional sparsity in the Copy task, where only tokens at fixed distances are attended to; content sparsity in the Select task, where only tokens with specific content are attended to; and associative sparsity in the Induce task, where only key-value pairs relevant to the query are attended to. These inherent sparsity patterns provide a theoretical foundation for designing more efficient attention mechanisms.

### v6

@@ Abstract
The increasing demand for long-context modeling in large language models (LLMs) is bottlenecked by the quadratic complexity of the standard self-attention mechanism. The community has proposed sparse attention to mitigate this issue. However, position-aware sparse attention methods rely on static sparse structures that lack adaptability to diverse query contexts, while content-aware sparse attention methods depend on heuristic key-value selection, hindering full differentiability. We introduce a trainable dynamic mask sparse attention mechanism, a method that merges the advantages of both position-aware and content-aware approaches. Dynamic Mask Attention (DMA) achieves this through three key innovations: First, it leverages value vector representations to generate content-aware dynamic masks, enabling the model to adaptively identify and attend to critical information. Second, it computes position-aware sparse weights in a hardware-friendly manner, efficiently skipping unnecessary computational regions. Finally, we demonstrate that the introduced dynamic mask and sparse weights do not obstruct gradients, supporting end-to-end training. We have validated the performance of DMA through comprehensive experiments. A large body of experimental evidence shows that DMA consistently holds a Pareto advantage over state-of-the-art sparse attention baselines in tasks including scaling laws, multi-query associative recall, standard benchmarks, and needle in a haystack tests, while also delivering up to a 10x overall speedup. These results highlight its ability to effectively balance model efficiency with long-context modeling capabilities. Our computational kernel code is now open-source at https://github.com/flash-algo/flash-sparse-attention to encourage further research and application by the community.
@@ 1 Introduction
Recent breakthroughs in large language models (LLMs) have yielded remarkable achievements in tasks requiring long-context reasoning [snell2024tts], such as deep reasoning [hf2025openr1], codebase generation [zhang2024codeagent], and multi-turn autonomous agents [park2023generative]. A key factor underpinning these successes is the ability to effectively model long-range dependencies, often spanning thousands of tokens [qwen32025, gemini2025, deepseekai2025deepseekr1incentivizingreasoningcapability]. However, the standard self-attention mechanism [vaswani2017attention] employed in Transformer architectures inherently exhibits quadratic computational complexity [zaheer2020big], which severely restricts scalability to longer sequences. Consequently, designing attention mechanisms that enhance computational efficiency without sacrificing modeling accuracy has become a critical research frontier for advancing the capabilities of LLMs.
Limitations of Existing Methods.
Current efficient attention strategies primarily leverage two types of sparsity: the position sparsity [martins2016softmax], which facilitates the efficient computation of essential query-key pairs, and the content sparsity [ge2023model], which enables the selective computation of relevant tokens. The first category includes methods such as sliding window attention [beltagy2020longformerlongdocumenttransformer], which employs static structures; multi-head latent attention [deepseekai2025deepseekv3technicalreport], which uses low-rank approximations; and native sparse attention [yuan2025nativesparseattentionhardwarealigned], which utilizes learnable compression weights. Although these approaches can achieve efficient long-context modeling, they often struggle to maintain their effectiveness. The second category encompasses KV cache eviction methods [li2024snapkv, zhang2023h2o, zhou2024llm]; block-wise KV cache selection strategies that dynamically choose cache blocks based on relevance predictions [gao2024seerattention, tang2024quest, xiao2024infllm]; and filtering methods that employ sampling [chen2024magicpig], clustering [liu2024clusterkv], or hashing [desai2024hashattention]. Despite their conceptual appeal, these techniques often fail to realize their theoretical speedups in practical deployments due to the overhead from dynamic computations or inaccurate sparsification decisions.
Key Challenges.
To overcome these limitations, an ideal sparse attention mechanism must simultaneously address two fundamental challenges: leveraging position-aware sparsity for essential computations [child2019generating] and exploiting content-aware sparsity for selective computation [dai2019transformerxl]. Meeting both requirements is crucial for achieving efficient and effective long-context reasoning and training in practice. However, existing methods still exhibit limitations, often facing a trade-off between efficiency and effectiveness. This dilemma highlights the urgent need for attention mechanisms that can preserve information integrity while achieving computational efficiency.
Figure 1: Workflow and Performance of Dynamic Mask Attention. Left: Overall workflow of DMA. The first step projects the input into $Q$, $K$, and $V$. The second step generates content-aware dynamic masks. The third step computes sparse weights. Black solid arrows indicate the forward computation path, while gray dashed arrows represent the backward computation path. Right: Relative performance comparison between full attention and DMA on benchmark tests. DMA achieves higher recall rates and significantly faster speeds while maintaining competitive accuracy.
@@
Table 1: Comparison of Different Attention Variants. Comparison of different Self-Attention mechanisms. $n$ denotes sequence length, $d_{h}$ represents head dimension, $w$ is window size, $d_{c}$ is compressed dimension, $B$ is compression block size, and $k$ is selection budget. Complexities focus on attention weight computation and memory requirements. Trainable indicates whether the sparsity pattern can be learned end-to-end.
Mechanism Comp. Complexity Mem. Complexity Sparsity Trainable
MHA $O(n^{2}d_{h})$ $O(n^{2})$ Static
SWA $O(nwd_{h})$ $O(nw)$ Position-aware
MLA $O(n^{2}d_{c})$ $O(n^{2})$ Low-rank Approx
NSA $O(n^{2}d_{c}/B+nkBd_{h}+nwd_{h})$ $O(n^{2}/B+nkB)$ Hybrid
H2O $O(nkd_{h})$ $O(nk)$ Content-aware
InfLLM $O(nkd_{h})$ $O(nk)$ Content-aware
Quest $O(nkd_{h})$ $O(nk)$ Content-aware
DAM $O(nkd_{h})$ $O(nk)$ Content-aware
DMA $O(nwd_{h})$ $O(nw)$ Content-Position Dual-aware
Figure 2: Dynamic Mask Attention Architecture. Left: Content-Aware Mask Computation. The mask computation part of dynamic mask attention. In the outer loop, the stride weight $\Delta$ and gate weight AA are loaded into high-speed SRAM, and in the inner loop, the zero-order hold method is used to loop through the $V$ blocks loaded into SRAM, sampling from it to generate content-aware $K$ masks. These masks are then causally broadcast to the length of $Q$ in HBM. Finally, in the outer loop, all mask blocks are concatenated to form the final content-aware sparse dynamic mask. Right: Position-Aware Weights Computation. The weight computation part of dynamic mask attention, where in the outer loop, the $K$ and $V$ blocks are looped and loaded into SRAM, and in the inner loop, the $Q$ blocks are accessed, loaded into SRAM, and the output of the attention weight computation is written back to HBM. If the current position of the $K$ block is designated as masked in the dynamic mask, the attention weight at that position is directly filled with 0, skipping the computation at that position, forming the final position-aware sparse attention weights.
Our Method.
In order to address these challenges and achieve efficient and effective sparse attention mechanisms, we ingeniously integrate the strengths of both strategies, attempting to strike a balance between the two, and propose Dynamic Mask Attention (DMA) to tackle the challenges of long-context modeling. As shown in Table 1, compared to other attention variants, DMA is a trainable content-position dual-aware sparse attention mechanism. As illustrated in Figure 1, it leverages two core innovations: generating dynamic masks using content-aware sparsity and computing sparse weights using position-aware sparsity, allowing the model to focus on relevant tokens while ignoring irrelevant ones. Furthermore, all computational operations are designed to be continuously differentiable, enabling end-to-end training of dynamic mask attention via gradient descent.
Kernel Design.
We implement a dedicated CUDA kernel that merges the memory efficiency of FlashAttention [dao2022flashattention] with DMA’s trainable sparsity, as illustrated in Figure 2, enabling hardware-level skipping of masked regions without incurring additional redundant computations. The kernel natively supports attention masks and biases with batch, head, and query broadcasting, ensuring flexible integration with diverse Transformer architectures. A block-level reduction determines the skip logic: tiles corresponding to all-zero masks bypass both computation and memory access, reducing the effective complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\cdot w)$ for $w\ll n$. The forward and backward passes share a unified skip logic, fetching $K/V$ tiles only when necessary, thereby maintaining an $\mathcal{O}(n)$ memory footprint without materializing the full attention matrix. The backward pass incorporates a complete gradient chain with fused bias gradients, rendering the entire pipeline fully differentiable for end-to-end training. To maximize throughput, we employ shared memory aliasing, pipelined prefetching, and coalesced memory accesses to minimize bandwidth pressure and improve hardware occupancy. These design choices allow DMA to sustain high performance on extremely long contexts, such as 128K+ tokens, while preserving accuracy comparable to dense attention baselines.
Contributions.
We make the following contributions: (i) We introduce Dynamic Mask Attention (DMA), a trainable, content- and position-aware sparse attention mechanism that decouples content-driven dynamic mask generation from position-aware sparse weight computation, reducing the effective time complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\cdot w)$ and the memory footprint to $\mathcal{O}(n\cdot w)$ for window size $w\ll n$. (ii) We develop a fully differentiable CUDA kernel that fuses FlashAttention-style tiling with hardware-efficient mask skipping. The kernel supports batch/head/query broadcasting for masks and biases, employs block-level skip logic to avoid unnecessary computation and memory traffic, and maintains an $\mathcal{O}(n)$ memory footprint without materializing the full attention weight matrix. (iii) We validate DMA across diverse application scenarios, showing consistent speedups over other optimized implementations, improved scaling laws [hoffmann2022empirical], higher accuracy on multi-query associative recall [arora2024zoology], competitive downstream benchmarks performance, and robust needle-in-a-haystack [gkamradt2023needleinahaystack] results.
2 Rethinking Sparse Attention
Since the advent of the Transformer [vaswani2017attention], the attention mechanism has become central to sequence modeling, yet its $O(n^{2})$ computational and memory complexity remains a bottleneck for processing long sequences. To overcome this limitation, researchers have explored sparse attention from two core perspectives: position-aware essential computation and content-aware selective computation. The former reduces computational load through predefined sparse patterns, while the latter dynamically determines the scope of computation based on input content. This section reviews the evolution of sparse attention, analyzes the strengths and weaknesses of existing methods, and provides context for our proposed Dynamic Mask Attention.
2.1 Position-Aware Essential Computation
To achieve hardware efficiency, early sparse attention methods predominantly employed fixed sparse patterns, aiming to simplify computation through structured sparsity.
Static Locality. Sliding Window Attention [beltagy2020longformerlongdocumenttransformer] confines computation to a local neighborhood for each token, reducing complexity to $O(n\cdot w)$, where $w$ is the window size. While simple and efficient, its fixed local window limits the model’s ability to capture long-range dependencies. This limitation is particularly pronounced in tasks requiring information integration across window boundaries, such as long-form question answering or code analysis.
Low-Rank Approximation. Methods like Multi-Head Latent Attention [deepseekai2025deepseekv3technicalreport] approximate the full attention matrix using low-rank decomposition to reduce computational and memory demands. While this approach is better at preserving global information than sliding window attention, it comes at the cost of precision loss due to information compression. Low-rank approximation can obscure fine-grained details crucial for specific tasks and, due to its global nature, cannot dynamically adjust its compression strategy based on context, lacking content adaptability.
Hardware-Aligned Sparsity. Work such as Native Sparse Attention [yuan2025nativesparseattentionhardwarealigned] designs regularized sparse patterns for modern accelerators, achieving high computational efficiency through hardware-friendly block-sparse structures. However, the core deficiency of such methods lies in their static nature. Fixed sparse patterns cannot adapt to the dynamic changes in input content, leading to potential misallocation of computational resources to non-critical regions while neglecting genuinely important information.
2.2 Content-Aware Selective Computation
As model capabilities have advanced, research has shifted towards content-aware selective computation, enabling models to learn autonomously where to focus their attention.
KV Eviction. Methods like H2O [zhang2023h2o] and SnapKV [li2024snapkv] save memory and computation by evicting "unimportant" tokens from the KV cache. These approaches typically rely on heuristics such as attention scores or access frequency to decide which tokens to retain. While effective in some scenarios, these heuristics can lead to erroneous eviction decisions, permanently losing critical information, especially in complex reasoning tasks that require long-distance backtracking.
Token Selection. Another class of methods actively selects a small subset of tokens for attention computation through techniques like sampling [chen2024magicpig], hashing [desai2024hashattention], or clustering [liu2024clusterkv]. While theoretically appealing, these methods face two major challenges in practice. First, discrete selection operations (such as sampling and hashing) are often non-differentiable, which impedes end-to-end training and prevents the model from learning optimal sparse patterns. Second, token-granular selection strategies disrupt memory access continuity, rendering them incompatible with modern efficient attention implementations like FlashAttention, leading to low hardware utilization and a decrease in both training and inference speed.
In summary, existing sparse attention mechanisms face an inherent trade-off between efficiency and effectiveness. Position-based methods, while efficient, lack flexibility and content awareness. Content-based methods, while more intelligent, are often limited by non-differentiable operations and inefficient hardware implementations. This dilemma highlights the urgent need for a new attention paradigm that can leverage structured sparsity for efficient computation while enabling content-aware selection through a trainable, dynamic mechanism.
@@ 3 Background
Figure 3: Sparsity in Language Modeling Tasks. The tasks of Copy, Select, and Induce are three essential tasks for language modeling. The Copy task requires maintaining a fixed distance between input elements and output elements, the Select task involves selectively remembering or ignoring certain elements based on the input, and the Induce task requires retrieving answers through associative recall based on context. Where the colored parts represent the tokens that the model needs to remember in the current time step of inference, the black parts represent the output tokens that the model needs to predict based on the input, and the white parts represent irrelevant tokens that can be filtered out.
@@ 3.1 Language Modeling Tasks
Sparsity in Language Modeling. As illustrated in Figure 3, long-context language modeling can be decomposed into three fundamental tasks: Copying [romero2021ckconv], Selecting [arjovsky2016unitary], and Inducing [olsson2022context]. The Copy task requires preserving fixed-distance relationships between input and output tokens. The Select task involves selectively retaining or discarding information based on its content. The Induce task necessitates retrieving information via associative recall from the context. Each of these tasks is characterized by a distinct sparsity pattern: the Copy task exhibits positional sparsity, attending only to tokens at fixed distances; the Select task demonstrates content sparsity, focusing on tokens with specific content; and the Induce task relies on associative sparsity, where attention is directed only to key-value pairs relevant to the query. These inherent sparsity patterns provide a strong theoretical foundation for designing more efficient attention mechanisms.